State and prove generalized commutative law in a commutative semigroup
1
Expert's answer
2019-07-29T09:27:34-0400

Generalized commutative law (GCL) :

let {i1, i2, ..., in} be the permutation of {1, 2, .., n}, then for elements a1, a2, ... an of the semigroup we have


"a_1*a_2*...*a_n=a_{i_1}*a_{i_2}*...*a_{i_n}."

Proof (using extended mathematical induction):

1) For n = 1 we have a1 = a1, for n = 2 a1*a2 = a2*a1 therefor GCL holds for n = 1 and n = 2.

2)Assume that GCL is true for all n < k + 1.

3)Prove that GCL is true for n = k + 1, that is


"a_1*a_2*...*a_{k+1}=a_{i_1}*a_{i_2}*...*a_{i_{k+1}}."

Let


"a_{i_m}=a_{k+1}"

then using generalized associative property of the semigroup


"a_{i_1}*a_{i_2}*...*a_{i_{m-1}}*a_{i_{m}}*a_{i_{m+1}}*...*a_{i_{k+1}}="

"(a_{i_1}*a_{i_2}*...*a_{i_{m-1}})*(a_{i_{m}}*(a_{i_{m+1}}*...*a_{i_{k+1}}))"


using 2)

"(a_{i_1}*a_{i_2}*...*a_{i_{m-1}})*(a_{i_{m}}*(a_{i_{m+1}}*...*a_{i_{k+1}}))="


"(a_{i_1}*a_{i_2}*...*a_{i_{m-1}})*((a_{i_{m+1}}*...*a_{i_{k+1}})*a_{i_{m}})"

using generalized associative property again


"(a_{i_1}*a_{i_2}*...*a_{i_{m-1}})*((a_{i_{m+1}}*...*a_{i_{k+1}})*a_{i_{m}})="

"(a_{i_1}*a_{i_2}*...*a_{i_{m-1}}*a_{i_{m+1}}*...*a_{i_{k+1}})*a_{i_{m}}"

(i1, i2,..., im-1, im+1,...,ik+1 is a permutation of 1, 2 , ..., k) using 2) for n = k


"(a_{i_1}*a_{i_2}*...*a_{i_{m-1}}*a_{i_{m+1}}*...*a_{i_{k+1}})="

"a_1*a_2*...*a_k"

therefor


"(a_{i_1}*a_{i_2}*...*a_{i_{m-1}}*a_{i_{m+1}}*...*a_{i_{k+1}})*a_{i_{m}}="

"a_1*a_2*...*a_k*a_{i_m}=a_1*a_2*...*a_k*a_{k+1}"

and


"a_{i_1}*a_{i_2}*...*a_{i_{k+1}}="

"a_{i_1}*a_{i_2}*...*a_{i_{m-1}}*a_{i_{m}}*a_{i_{m+1}}*...*a_{i_{k+1}}="

"a_1*a_2*...*a_k*a_{k+1}"

So we have proved GCL using extended mathematical induction.